8. Rocephin O: 750 mg IM daily x 2 days beginning today. A: 1gm/ml Give: ml
Answer #1

Ordered dose=750 mg

Dose in hand= 1000 gm= 1gm

1 ml contains 1gm medicine

Medicine administered in ml= 750÷1000=3/4=0.75 ml